### v1

# Neural Beam Field for Spatial Beam RSRP Prediction

Abstract
Accurately predicting beam-level reference signal received power (RSRP) is essential for beam management in dense multi-user wireless networks, yet challenging due to high measurement overhead and fast channel variations. This paper proposes Neural Beam Field (NBF), a hybrid neural-physical framework for efficient and interpretable spatial beam RSRP prediction. Central to our approach is the introduction of the Multi-path Conditional Power Profile (MCPP), which bridges site-specific multipath propagation with antenna/beam configurations via closed-form analytical modeling. We adopt a decoupled “blackbox-whitebox” design: a Transformer-based deep neural network (DNN) learns the MCPP from sparse user measurements and positions, while a physics-inspired module analytically infers beam RSRP statistics. To improve convergence and adaptivity, we further introduce a Pretrain-and-Calibrate (PaC) strategy that leverages ray-tracing priors and on-site calibration using RSRP data. Extensive simulations results demonstrate that NBF significantly outperforms conventional table-based channel knowledge maps (CKMs) and pure blackbox DNNs in prediction accuracy, training efficiency, and generalization, while maintaining a compact model size. The proposed framework offers a scalable and physically grounded solution for intelligent beam management in next-generation dense wireless networks.
Index Terms:
Neural Beam Field, Multipath Conditional Power Profile, Channel Knowledge Map, Beam‐level RSRP, Transformer
I Introduction
With the continuous densification of mobile communication networks, network operators face unprecedented challenges in maintaining scalable, highly spectral-efficient and reliable service coverage. Although massive Multiple-Input Multiple-Output (MIMO) and millimeter-wave technologies promise significant improvements in capacity and user throughput, the practical deployment of dense wireless networks is significantly constrained by the difficulty of timely acquiring Channel State Information (CSI) under realistic multi-user scenarios[1]. Traditionally, accurate CSI acquisition involves comprehensive measurements at high spatial-temporal-spectral resolutions, which inevitably leads to prohibitive overheads and excessive latency, severely limiting network scalability and real-time applicability [2, 3].
Traditional channel estimation for MIMO beamforming mainly relies on two methods, i.e., pilot-based CSI estimation and beam scanning. The former method provides full CSI but incurs high overhead and latency, with the pilots given in the fifth-generation (5G) standards in the form of downlink CSI-RS or uplink SRS signals[2, 3]. In comparison, the beam scanning method selects beam pairs from a codebook to estimate Reference Signal Received Power (RSRP) or Signal-to-Interference-plus-Noise Ratio (SINR) for user-specific beam pair links (BPLs). To alleviate the burden of CSI acquisition, reduced-dimensional channel indicators such as RSRP or SINR have been widely adopted in practical network deployments [4, 5]. Nevertheless, even these simplified RSRP-based measurements can hardly satisfy real-time requirements due to varying propagation conditions and user mobility. Moreover, dense network conditions exaggerate the complexity, as complete measurement of RSRP for all user-beam combinations often exceeds channel coherence time, rendering CSI outdated before it can be utilized for resource allocation or beam scheduling [6]. Moreover, the beam-specific RSRP is highly directional and thus relies on the site-specific blockages/shadowing/multi-path conditions. These critical issues demand innovative approaches capable of effectively predicting spatial RSRP fields under diverse beamforming configurations and environmental conditions.
Recent developments propose various spatial interpolation and estimation techniques to overcome CSI acquisition challenges by constructing radio map[7] or channel knowledge map (CKM)[8] from limited spatial samples. Early methods commonly assume spatial smoothness of received power, employing techniques such as Kriging [9], thin-plate spline interpolation [10], kernel methods [11], and Gaussian radial basis functions [12]. Alternative methods, including sparse representation [13] and low-rank matrix or tensor completion [14], have also validated theoretical feasibility under structural priors like sparsity or low-rankness. A new scatterer-based channel model is also proposed in [15] for long-scale channel gain map estimation. However, these methods might suffer from notable performance degradation in complex urban environments characterized by dynamic multipath propagation, blockage effects, and positioning uncertainties[7, 8].
In contrast, deep learning methods have shown substantial promise in capturing intricate spatial-temporal correlations and providing robust predictions, offering new opportunities for accurate prediction of RSRP [16, 17, 18]. Inspired by Neural Radiance Fields (NeRF) [19] from computer vision, recent works further explore deep neural network (DNN)-based spatial field estimation methods integrated with physical propagation mechanisms. Notably, Neural RF Radiation Fields (NeRF2) [16] predict signal features at arbitrary locations from sparse measurements, while Neural Wireless Radiation Fields (NeWRF) explicitly incorporate electromagnetic propagation physics into neural frameworks to enhance modeling accuracy in heterogeneous environments [17]. Additionally, WiNeRT [18] leverages DNNs to model intricate interactions between rays and reflective surfaces, capturing complex multipath propagation dynamics. Nevertheless, these approaches have yet to explicitly incorporate beam-specific RSRP measurements in practical 5G deployments.
Motivated by these limitations, we introduce the Neural Beam Field (NBF), a novel DNN-based method tailored for predicting spatial beam RSRP statistics, explicitly incorporating directional beamforming and multipath propagation physics. Our main contributions are summarized as follows:
  • We introduce the new notion of Multi-path Conditional Power Profile (MCPP), which serves as an essential intermediary between physical wave propagation in a site-specific environment and the add-on antenna panel/beam configurations. Based on MCPP, we further derive analytical formulas for beam RSRP statistics, which facilitate easier learning and generalization under diverse settings.
  • We propose a decoupled “blackbox-whitebox” approach to construct NBF, by 1) learning the site-specific/irregular MCPPs using a tailored design of Transformer-based DNN, and 2) inferring the beam RSRP statistics based on analytical formulas. This “whitebox” analytical modeling combined with “blackbox” neural representations results in an interpretable and accurate predictive framework.
  • To achieve better convergence and optimality for the highly nonlinear mapping task in NBF, we further propose a Pretrain-and-Calibrate (PaC) strategy in cases when prior (but not necessarily exact) information about MCPPs are available. Leveraging on tailored MCPP loss functions, the PaC strategy helps tame the NBF into a desired locally optimal point, which could then adapt to unknown environmental factors by on-site RSRP measurements.
Numerical results validate the analytical formulas through extensive Monte Carlo (MC) simulations. Moreover, it is shown that the proposed approach significantly outperforms conventional table-based CKMs and pure black-box neural models in terms of prediction accuracy, storage efficiency, and generalization capability. Our proposed NBF method for spatial beam RSRP prediction holds potential in intelligent beam management and user scheduling in dense wireless networks.
Notations: Symbols for vectors (lower case) and matrices (upper case) are in boldface. $\otimes$ ($\odot$) represents the Kronecker (Hadamard) product. $(\cdot)^{T}$ and $(\cdot)^{*}$ represent the transpose and conjugate, respectively. Re($\cdot$) takes the real part of a complex number while $\angle\cdot$ takes the angle of a complex number.
II System Model
Figure 1: Three‐sector urban macrocell scenario with site-specific blockages and multipath propagation conditions.
As shown in Fig. 1, we consider downlink communications in a typical three‐sector urban macrocell scenario with site-specific blockages and multipath propagation conditions, where the service region $\mathcal{A}$ is partitioned into regular grids with spacing $d$ meters (m). For simplicity, the user equipments (UEs) are assumed to be equipped with a single isotropic antenna at height $h_{\textrm{UE}}$ m. The base station (BS) is located at a known position with height $h_{\textrm{BS}}$ m, and employs three sector antenna panels each configured as a uniform planar array (UPA) of size $N_{\text{tx}}\triangleq N_{\text{tx,y}}\times N_{\text{tx,z}}$. Each antenna element is assumed to exhibit a certain element radiation pattern (ERP), e.g., the directional antenna model as in 3GPP TR 38.901[20], denoted by $G\left(\Pi_{\text{tx}}\right)$, which represents the power pattern as seen from the azimuth and elevation angles $\Pi_{\text{tx}}\triangleq\left(\theta_{\text{tx}},\varphi_{\text{tx}}\right)$ in the global coordinate system (GCS), whose origin co-locates with the reference BS antenna element.
II-A Antenna Panel Rotation and Array Response Vector
Each antenna panel $i\in\mathcal{I}=\{1,2,3\}$ has counterclockwise rotation angles $\rho_{\text{x},i}$, $\rho_{\text{y},i}$ and $\rho_{\text{z},i}$ about the $x$, $y$, and $z$ axes, respectively, based from the default Y-O-Z plane in the GCS. Let $\mathbf{n}_{0}$, $\mathbf{h}_{0}$ and $\mathbf{v}_{0}$ denote the GCS orthonormal basis vectors corresponding to the normal, horizontal and vertical directions, respectively. For a given panel, when the array is rotated by angles $\rho_{\text{x}}$, $\rho_{\text{y}}$ and $\rho_{\text{z}}$ about the $x$, $y$, and $z$ axes in sequence, the resulting composite rotation matrix is given by [20]
$\mathbf{R}=\mathbf{R}_{\text{z}}(\rho_{\text{z}})\,\mathbf{R}_{\text{y}}(\rho_{\text{y}})\,\mathbf{R}_{\text{x}}(\rho_{\text{x}}).$ (1)
As a result, the basis vectors after rotation are given by
$\mathbf{n}=\mathbf{R}\,\mathbf{n}_{0},\quad\mathbf{h}=\mathbf{R}\,\mathbf{h}_{0},\quad\mathbf{v}=\mathbf{R}\,\mathbf{v}_{0},$ (2)
which form the orthonormal basis for the panel’s local coordinate system (LCS), as shown in Fig. 1.
Denote $\boldsymbol{u}_{\text{tx}}\triangleq[\cos\varphi_{\text{tx}}\sin\theta_{\text{tx}},\sin\varphi_{\text{tx}}\sin\theta_{\text{tx}},\cos\theta_{\text{tx}}]^{T}$ as the 3-dimensional (D) unit direction vector of a given wave propagation path w.r.t. the origin, with $\Pi_{\text{tx}}\triangleq(\varphi_{\text{tx}},\theta_{\text{tx}})$ denoting the corresponding (horizontal, vertical) direction of departure (DOD) angle in the GCS. Consider one typical antenna panel which locates in the Y-O-Z plane in the GCS. The components of its array response vector (ARV) along the $y\text{-}$ and $z\text{-}$axis are given by
$\boldsymbol{a}_{\text{tx,y}}(\Pi_{\text{tx}})\triangleq[1,\cdots,e^{j\frac{2\pi f_{\text{c}}}{c}(\boldsymbol{r}_{N_{\text{tx,y}}}^{T}\boldsymbol{u}_{\text{tx}})}]^{T}\in\mathbb{C}^{N_{\text{tx,y}}\times 1},$ (3)
$\boldsymbol{a}_{\text{tx,z}}(\Pi_{\text{tx}})\triangleq[1,\cdots,e^{j\frac{2\pi f_{\text{c}}}{c}(\boldsymbol{r}_{N_{\text{tx,z}}}^{T}\boldsymbol{u}_{\text{tx}})}]^{T}\in\mathbb{C}^{N_{\text{tx,z}}\times 1},$ (4)
where $\boldsymbol{r}_{n_{\text{tx,y}}}\triangleq(n_{\text{tx,y}}-1)d_{\text{y}}\mathbf{h}$ and $\boldsymbol{r}_{n_{\text{tx,z}}}\triangleq(n_{\text{tx,z}}-1)d_{\text{z}}\mathbf{v}$ denote the position vectors of the $(n_{\text{tx,y}},n_{\text{tx,z}})$-th element along the $y\text{-}$ and $z\text{-}$axis, with $d_{\text{y}}$ and $d_{\text{z}}$ being the element spacing, respectively; $c$ denotes the speed of light and $f_{\text{c}}$ denotes the carrier frequency. Then, the ARV is given by $\boldsymbol{a}_{\text{tx}}(\Pi_{\text{tx}})\triangleq\boldsymbol{a}_{\text{tx,y}}(\Pi_{\text{tx}})\otimes\boldsymbol{a}_{\text{tx,z}}(\Pi_{\text{tx}})\in\mathbb{C}^{N_{\text{tx}}\times 1}$. The ARV for other panels with different rotations can be expressed in similar manners.
II-B Channel Model
For the pure propagation channel (assuming an isotropic antenna at both sides) between the BS and the target UE, assume that there exists $L$ significant propagation paths. Each path $l\in\mathcal{L}\triangleq\{1,\cdots,L\}$ is associated with unit DOD vector $\boldsymbol{u}_{\text{tx},l}$ and unit direction of arrival (DOA) vector $\boldsymbol{u}_{\text{rx},l}\triangleq[\cos\varphi_{\text{rx},l}\sin\theta_{\text{rx},l},\sin\varphi_{\text{rx},l}\sin\theta_{\text{rx},l},\cos\theta_{\text{rx},l}]^{T}$, with $\Pi_{\text{rx},l}\triangleq(\varphi_{\text{rx},l},\theta_{\text{rx},l})$ denoting the corresponding (horizontal, vertical) DOA angle w.r.t. the UE position. As a result, the BS-UE channel for the $l$-th path is given by
$\boldsymbol{g}_{l}\triangleq\textstyle\sqrt{G(\Pi_{\text{tx},l})}\alpha_{l}e^{j\Phi_{l}}e^{-j2\pi f_{\text{c}}\tau_{l}}\boldsymbol{a}_{\text{tx}}(\Pi_{\text{tx},l}),$ (5)
where $\alpha_{l}$, $\Phi_{l}$ and $\tau_{l}$ denote the amplitude, phase and propagation delay of the $l$-th path, respectively. Note that for a given UE position in a quasi-static environment, the per-path geometric parameters including $\Pi_{\text{tx},l}$, $\Pi_{\text{rx},l}$, $\alpha_{l}$ and $\tau_{l}$ mainly depend on the environment geometry and could be obtained from ray tracing or site survey and stored in the form of CKMs. On the other hand, the channel phase $\Phi_{l}$ is affected by the micro user movements/environment changes, which is thus treated as a random variable.
Therefore, we introduce the notion of Single-path Conditional Power (SCP) for each path $l\in\mathcal{L}$, denoted as
$\mathrm{SCP}_{l}=\alpha_{l}^{2}\triangleq p_{l}\bigl{(}\boldsymbol{u}_{\text{tx},l},\,\boldsymbol{u}_{\text{rx},l},\,\tau_{l}\bigr{)},$ (6)
where $(\boldsymbol{u}_{\text{tx},l},\,\boldsymbol{u}_{\text{rx},l},\,\tau_{l}\bigr{)}$ collectively represents the seven-dimensional (7-D) path condition. Note that we choose the unit direction vectors $\boldsymbol{u}_{\text{tx},l}$ and $\boldsymbol{u}_{\text{rx},l}$ instead of the angles $\Pi_{\text{tx},l}$ and $\Pi_{\text{rx},l}$ for better representation and easier learning of periodic angle features. Furthermore, we introduce the new notion of Multi-path Conditional Power Profile (MCPP) for all paths in $\mathcal{L}$, denoted as
$\mathrm{MCPP}\triangleq\{\mathrm{SCP}_{1},\,\mathrm{SCP}_{2},\,\dots,\,\mathrm{SCP}_{L}\},$ (7)
which is a set of $L$ power entries in the 7-D conditional space.
Remark: The MCPP for the pure propagation channel serves as an essential intermediary between physical wave propagation in a site-specific environment and the add-on antenna/beam configurations, which jointly determine the beam RSRP statistics. Therefore, we propose a decoupled “blackbox-whitebox” approach by 1) learning the site-specific/irregular MCPPs using DNNs, and 2) inferring the beam RSRP statistics based on our derived analytical formulas in the following.
II-C Beam RSRP Statistics under Given Multipath Conditions
II-C1 RSRP under Given Beamforming Scheme
For the purpose of exposition, here we consider analog beamforming with discrete Fourier transform (DFT) codebooks. [^1] Note that our proposed method is general and can be extended to other beamforming schemes such as hybrid digital/analog beamforming. Specifically, for the typical UPA panel in the Y-O-Z plane, the components of its beamforming vector along the $y\text{-}$ and $z\text{-}$axis are given by
$\boldsymbol{w}_{\text{tx,y}}(\xi_{\text{tx,y}})\triangleq[1,e^{j\xi_{\text{tx,y}}},\cdots,e^{j(N_{\text{tx,y}}-1)\xi_{\text{tx,y}}}]^{T}\in\mathbb{C}^{N_{\text{tx,y}}\times 1},$ (8)
$\boldsymbol{w}_{\text{tx,z}}(\xi_{\text{tx,z}})\triangleq[1,e^{j\xi_{\text{tx,z}}},\cdots,e^{j(N_{\text{tx,z}}-1)\xi_{\text{tx,z}}}]^{T}\in\mathbb{C}^{N_{\text{tx,z}}\times 1},$ (9)
where $\xi_{\text{tx,y}}$, $\xi_{\text{tx,z}}\in[-\pi,\pi)$ denote the spatial frequencies along the $y\text{-}$ and $z\text{-}$axis, respectively. Then, the overall DFT-type beamforming vector with normalized power is given by
$\boldsymbol{w}_{\text{tx}}(\xi_{\text{tx,y}},\xi_{\text{tx,z}})\triangleq\frac{1}{\sqrt{N_{\text{tx}}}}\boldsymbol{w}_{\text{tx,y}}(\xi_{\text{tx,y}})\otimes\boldsymbol{w}_{\text{tx,z}}(\xi_{\text{tx,z}}).$ (10)
In the downlink, the received signal is then given by
$y\triangleq\textstyle\sum\nolimits_{l\in\mathcal{L}}\boldsymbol{g}_{l}^{T}\,\mathbf{w}_{\text{tx}}\,s\;+\;n,$ (11)
where $s$ denotes the transmit symbol with power $P_{\text{t}}$ Watt (W) and $n$ denotes the additive white Gaussian noise. The beam RSRP normalized to $P_{\text{t}}$ is then given by
$P_{\text{r}}(\mathrm{MCPP},\mathbf{w}_{\text{tx}})\triangleq|\textstyle\sum\nolimits_{l\in\mathcal{L}}\boldsymbol{g}_{l}^{T}\,\mathbf{w}_{\text{tx}}|^{2},$ (12)
with $P_{\text{r,dB}}$ denoting the corresponding normalized RSRP in dB.
II-C2 Beam RSRP Statistics under Given MCPP
Here we derive the beam RSRP statistics under given MCPP. For simplicity, we first consider the narrowband scenario while leaving the wideband case for future extension. The results are summarized in the following proposition.
Proposition 1.
For independent random channel phases $\Phi_{l}$, $l\in\mathcal{L}$ uniformly distributed in $0,2\pi)$, the mean and variance of the normalized beam RSRP in (12) are given by
$\mu_{\mathrm{RSRP}}(\mathrm{MCPP},\mathbf{w}_{\text{tx}})=\textstyle\sum\nolimits_{l\in\mathcal{L}}\gamma_{l},$ (13)
$\sigma_{\mathrm{RSRP}}^{2}(\mathrm{MCPP},\mathbf{w}_{\text{tx}})=\left(\textstyle\sum\nolimits_{l\in\mathcal{L}}\gamma_{l}\right)^{2}-\textstyle\sum\nolimits_{l\in\mathcal{L}}\gamma_{l}^{2},$ (14)
where $\gamma_{l}\triangleq G(\Pi_{\text{tx},l})p_{l}|\Delta_{l}|^{2}$ denotes the average power gain, and $|\Delta_{l}|^{2}=\frac{1}{N_{\text{tx}}}\big{|}S_{N_{\text{tx,y}}}(\zeta_{\text{tx,y},l}+\xi_{\text{tx,y}})\big{|}^{2}\big{|}S_{N_{\text{tx,z}}}(\zeta_{\text{tx,z},l}+\xi_{\text{tx,z}})\big{|}^{2}$ denotes the power gain of the array factor $\Delta_{l}\triangleq\boldsymbol{a}_{\text{tx}}^{T}(\Pi_{\text{tx},l})\,\mathbf{w}_{\text{tx}}$, for path $l\in\mathcal{L}$, with $\zeta_{\text{tx,y},l}\triangleq\frac{2\pi f_{\text{c}}d_{\text{y}}}{c}\mathbf{h}^{T}\boldsymbol{u}_{\text{tx}}(\Pi_{\text{tx},l})$, $\zeta_{\text{tx,z},l}\triangleq\frac{2\pi f_{\text{c}}d_{\text{z}}}{c}\mathbf{v}^{T}\boldsymbol{u}_{\text{tx}}(\Pi_{\text{tx},l})$ and $S_{N}(\cdot)$ given by (19).
[…]
II-D Practical Collection/Synthetic Generation of RSRP Data
In current 5G standards, beam RSRP could be obtained from pilot signals such as synchronization signal blocks (SSBs) and CSI-reference signal (CSI‐RS). These measurement results could be combined with UE location information to construct a CKM for spatial beam RSRP prediction. In practical scenarios, however, the RSRP measurements might be performed by different UEs at different time and locations with possible positioning errors, which requires effective ways for data aggregation. This paper proposes a grid‐based weighted aggregation method for obtaining stable RSRP statistics. Specifically, the RSRP measurements are first mapped to neighboring grid anchors based on the estimated UE position and its confidence interval, with normalized sample weights based on, e.g., inverse distance weighting (IDW). Furthermore, for timely update of the CKM, a temporal weight decay mechanism could be introduced by gradually decaying the sample weights over time. As a result, each grid anchor absorbs spatial/temporal RSRP measurements and accumulates stable RSRP statistics.
On the other hand, as a preliminary study, we also provide ways for synthetic generation of RSRP data. In compliance with the 3GPP TR38.901 hybrid channel model[20], we employ SionnaRT[21] to generate site-specific MCPPs for a given city region based on ray tracing, while utilizing QuaDRiGa[22] to generate spatially consistent random perturbations for the MCPPs to simulate unknown environmental factors. The final MCPP for RSRP generation at each location is then obtained as the weighted sum of the above deterministic and random MCPP components.
III Design and Calibration of Neural Beam Field
The objective of NBF is to predict the beam RSRP statistics under any given UE location $\mathbf{x}$ and beam configuration $\mathbf{w}_{\text{tx}}$. Based on the spatially distributed RSRP measurements, conventional table-based CKMs could be employed by storing the beam RSRP statistics in discrete map grids and performing interpolation in under-sampled locations/beams. However, it typically incurs large storage overhead to store the complete set of CKMs under various configuration conditions, while remaining challenging for accurate spatial interpolation/nonlinear inference. On the other hand, DNN-based blackbox models could be employed to learn the complicated nonlinear mappings between UE positions/beam configurations and the corresponding RSRP statistics. Nevertheless, pure blackbox models lack physical interpretation and typically require sufficiently sampled datasets from the high dimensional input space, thereby often associated with prolonged training time and difficulty of generalization beyond the training dataset.
In this section, we propose a decoupled “blackbox-whitebox” approach to construct the NBF, by 1) designing a tailored DNN based on Transformer[23] to learn the site-specific/irregular MCPPs, and 2) inferring the beam RSRP statistics based on our derived analytical formulas in Proposition 1. Such an approach leverages on the essential role of MCPP in amalgamating propagation channels and beam configurations, which thus enables easier learning and generalization.
Figure 2: NBF Network Architecture.
III-A Neural Beam Field Design
The overall NBF design is illustrated in Fig. 2, which consists of 1) a Transformer-based blackbox model $f_{\vartheta}(\mathbf{x})$ that predicts the MCPP at a given UE location $\mathbf{x}$, and 2) a whitebox model that maps to beam RSRP statistics based on Proposition 1. Combined together, the NBF function is given by
$\mathrm{NBF}(\mathbf{x},\mathbf{w}_{\text{tx}})\triangleq\big{\{}\mu_{\mathrm{RSRP}}\big{(}f_{\vartheta}(\mathbf{x}),\mathbf{w}_{\text{tx}}\big{)},\,\sigma_{\mathrm{RSRP}}^{2}\big{(}f_{\vartheta}(\mathbf{x}),\mathbf{w}_{\text{tx}}\big{)}\big{\}}.$ (15)
In the following, we focus on the DNN design.
Firstly, for the input of UE location $\mathbf{x}$, we adopt the random Fourier position embedding[24] to obtain a 256-D vector $\mathbf{X}_{\text{UE}}\in\mathbb{R}^{1\times 256}$, which can then be treated as a token for subsequent Transformer blocks. In addition, in order to output the MCPP which consists of $L$ SCPs, we also place $L$ learnable target tokens $\mathbf{X}_{\text{target}}\in\mathbb{R}^{L\times 256}$ at the input. Secondly, the concatenated input tokens $\mathbf{X}\triangleq[\mathbf{X}_{\text{UE}};\mathbf{X}_{\text{target}}]$ are processed by the Transformer encoder, comprising $N$ blocks of multi-head self-attention (MSA), MLP, layer normalization (LN), and residual connections[23]. Notably, the learnable target tokens attend to the UE location token along the blocks to extract key information related to location-specific MCPPs. Finally, each of the target tokens processed by Transformer is fed to two MLPs, i.e., 1) a two-layer regression head that produces 8 features corresponding to one SCP in (6), thereby composing the predicted MCPP $\hat{\mathbf{Y}}=f_{\vartheta}(\mathbf{x})$; and 2) a one-layer classification head with output $a_{l}\in\{0,1\}$ that tells the existence of path $l$ (in cases with fewer than $L$ paths).
III-B End-to-End Learning Based on RSRP Measurements
Based on the above NBF design, we then train the DNN such that the end-to-end (blackbox-whitebox) prediction of beam RSRP statistics matches as close as possible to the corresponding labels from measurement data. Regression loss functions such as the smooth L1 loss or mean absolute error (MAE) loss could be employed.
III-C MCPP-Prior Aided Pretraining and On-Site Calibration
Due to the highly nonlinear mapping task in NBF, the training process might trap in undesired local points. To achieve accelerated convergence and better optimality, we further propose a Pretrain-and-Calibrate (PaC) strategy in cases when prior (but not necessarily exact) information about MCPPs are available, e.g., through ray tracing in an environment with known geometry and materials.
In the pretraining phase, MCPP labels in the form of (7) are available, denoted as $\bar{\mathbf{Y}}$. We aim to minimize the difference between the predicted MCPP $\hat{\mathbf{Y}}$ and ground-truth MCPP $\bar{\mathbf{Y}}$. However, two issues exist, i.e., 1) some labels $\bar{\mathbf{Y}}$ might have paths fewer or more than $L$, and 2) the permutation-invariant path indices require per-path matching between $\hat{\mathbf{Y}}$ and $\bar{\mathbf{Y}}$. Regarding the first issue, note that $L$ could be chosen as the typical (maximum) number of significant paths under given environment type and carrier frequency. In some outlier locations with more than $L$ paths, clustering methods such as weighted K-means could be employed to merge them into $L$ significant paths. On the other hand, for locations with fewer than $L$ paths, we utilize the path existence prediction $a_{l}$ along with binary cross-entropy (BCE) loss to learn the situation. Regarding the second issue, we employ the Hungarian algorithm for one-to-one path matching. The above processing thus allows us to capture the difference between $\hat{\mathbf{Y}}$ and $\bar{\mathbf{Y}}$, using regression loss functions such as MAE. The overall loss function is then taken as the weighted sum of the MCPP regression loss and path existence BCE loss, whereby the details are omitted for brevity.
Finally, in the calibration phase, the on-site RSRP measurements could be utilized to fine-tune the NBF based on the process in Section III-B. In summary, the PaC strategy helps tame the NBF into a desired locally optimal point, which could then adapt quickly to other unknown environmental factors implicitly manifested in the on-site RSRP measurements.
IV Numerical Results
This section presents numerical results that validate the analytical formulas and evaluate the effectiveness of NBF. First, we verify the analytical formulas from Proposition 1 through extensive MC simulations. Second, we assess the performance of NBF by comparing it with two benchmark schemes, including a pure blackbox model based on MLP, and a table-based CKM with fixed interpolation rules using IDW. The MLP baseline performs end-to-end regression of user performance, using normalized user coordinates and sin-cos encoded DFT beam spatial frequencies as inputs. Its model size is slightly larger than that of the NBF to ensure a fair comparison of model architectures. The IDW baseline interpolates the MCPP at a target location from nearby samples and plugs the estimate into the white-box analytical expression to obtain user performance. The following parameters are used unless otherwise specified: $h_{\text{BS}}=20$ m, $h_{\text{UE}}=1.5$ m, $L=10$, $f_{c}=3.5$ GHz, antenna panel dimensions $N_{\mathrm{tx,y}}=8$ and $N_{\mathrm{tx,z}}=4$, a square area with side length 256 m, a panel down-tilt angle of $15^{\circ}$, peak antenna element gain of 8 dBi, and half-wavelength inter-element spacing.
IV-A Validation of Analytical Formulas
To assess the validity of the derived theoretical model under realistic channel conditions, we simulate an outdoor scenario using the QuaDRiGa toolbox. For each beam and receiver location, we compute the theoretical mean and standard deviation of the RSRP under given MCPPs. In the MC simulations, we average over $N_{\mathrm{mc}}=200$ independent channel realizations, apply the DFT beamforming weights in each realization, and compute the empirical means and standard deviation of the instantaneous received power.
To validate the accuracy of our theoretical model, we first compare the mean and standard deviation of the beam RSRP predicted by our analytical formulas with those obtained from the MC simulations. For a given location, as shown in Fig. 3, the results show good agreement across different beam configurations. Furthermore, we compare the spatial distribution of the mean RSRP over the entire map, as depicted in Fig. 4, which confirms the consistency between the theoretical predictions and simulation results.
Figure 3: (a) Mean and (b) standard deviation of beam RSRP obtained by MC simulations and formulas under different beam configurations.
Figure 4: Comparions of spatial mean RSRP heatmap for a representative beam over the entire map region, obtained by either MC or formulas.
IV-B Effectiveness of NBF
For NBF training, we collect ray-traced user channels and RSRP samples on a site-specific map region. Each sample consists of the UE position, antenna panel and beam configurations, together with the associated performance label. We use an $80/20$ split of dataset for training and validation, and optimize the Smooth-$L_{1}$ loss on the mean RSRP using the Adam optimizer and OneCycleLR learning rate schedule.
Fig. 5 compares the training and validation loss curves for three approaches, i.e., 1) NBF with end-to-end learning from RSRP data, 2) NBF fine-tuning after MCPP-prior aided pretraining, and 3) an MLP baseline, also with end-to-end learning from RSRP data. First, it can be seen that NBF outperforms the MLP baseline in terms of convergence speed and prediction accuracy, both under end-to-end training. Second, the MCPP-prior aided pretraining helps tame the NBF into better initial points and further reduce both the training and validation losses, thus improving prediction accuracy and generalization.
Next, we evaluate the effectiveness of NBF through an illustrative example in Fig. 6, which presents spatial mean RSRP heatmaps for a representative beam under ground truth, pretrained NBF, and MLP. It can be seen that the NBF prediction closely follows the ground-truth distribution and captures fine spatial details, while the MLP output appears overly smooth and misses local details.
Figure 5: Training and validation loss curves for end-to-end NBF training (left), NBF fine-tuned after MCPP prior pretraining (middle), and an MLP baseline trained end-to-end (right).
Figure 6: Spatial mean RSRP heatmap comparison for a representative beam: ground truth (left), NBF prediction (center), and MLP prediction (right).
In terms of quantitative metrics, we evaluate the MAE between the predicted and ground-truth mean RSRPs (normalized to transmit power $P_{\text{t}}$) in dB scale, as well as the storage requirement in mega-bytes (MB) for CKM tables or DNN parameters, under three different schemes. Our NBF architecture under end-to-end training achieves MAE of $3.547$ dB while requiring only $8.21$ MB of storage. The MLP baseline recordes MAE of $4.105$ dB and requires $10.57$ MB of storage. These results show that NBF achieves higher prediction accuracy with lower storage requirement. When prior information about MCPP is available, we apply the MCPP prior to both the NBF pretraining and the IDW baseline, both of which utilize our whitebox analytical formulas to map MCPP into beam RSRP statistics. In this case, the MCPP-prior aided pretraining helps reduce its MAE from $3.547$ dB to $2.416$ dB, while IDW yields MAE of $3.328$ dB and consumes $27.86$ MB of storage. As a result, NBF outperforms IDW in fitting quality and storage efficiency.
V Conclusions
We proposed the NBF, a hybrid neural-physical framework for spatial beam RSRP prediction, based on MCPP as an essential intermediary to amalgamate propagation channels and antenna/beam configurations. The derived closed-form RSRP statistics were validated through Monte Carlo simulations under ray-traced multipath environments. Compared with conventional table-based CKMs with fixed interpolation rules (e.g., IDW), and pure blackbox models such as MLP, our proposed NBF achieved superior prediction accuracy, better generalization, and higher storage efficiency. The Pretrain-and-Calibrate strategy further improved convergence and adaptability. These results highlight NBF as an accurate, interpretable, and efficient solution for beam management in dense wireless networks. Future work may extend to wideband channels and hybrid beamforming schemes.
Appendix A Proof of Proposition 1
Based on (12), the normalized beam RSRP is given by
$P_{\text{r}}=|\textstyle\sum\nolimits_{l\in\mathcal{L}}\boldsymbol{g}_{l}^{T}\,\mathbf{w}_{\text{tx}}|^{2}=|\textstyle\sum\nolimits_{l\in\mathcal{L}}A_{l}e^{j\Phi_{l}}|^{2}$
$=\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}}A_{l}A_{m}^{*}e^{j(\Phi_{l}-\Phi_{m})}$
$\stackrel{{\scriptstyle(a)}}{{=}}\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}}\text{Re}\left\{A_{l}A_{m}^{*}e^{j(\Phi_{l}-\Phi_{m})}\right\}$
$=\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}}P_{l,m}\cos{\Psi_{l,m}},$ (16)
where $A_{l}\triangleq\sqrt{G(\Pi_{\text{tx},l})}\,\alpha_{l}\,e^{-j2\pi f_{\text{c}}\tau_{l}}\Delta_{l}$, $\Delta_{l}\triangleq\boldsymbol{a}_{\text{tx}}^{T}(\Pi_{\text{tx},l})\,\mathbf{w}_{\text{tx}}$, $P_{l,m}\triangleq\sqrt{G(\Pi_{\text{tx},l})G(\Pi_{\text{tx},m})}\,\alpha_{l}\alpha_{m}\left|\Delta_{l}\right|\left|\Delta_{m}\right|$, $\Psi_{l,m}\triangleq\Phi_{l}-\Phi_{m}+\angle\Delta_{l}-\angle\Delta_{m}-2\pi f_{\text{c}}(\tau_{l}-\tau_{m})$, and the equation $(a)$ is due to $\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}}A_{l}A_{m}^{*}=\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}}\text{Re}(A_{l}A_{m}^{*})$ in general. Note that the random phases $\Phi_{l},l\in\mathcal{L}$ appear in the term $\Psi_{l,m}$. For the case with $l=m$, we have $\Psi_{l,m}=0$ and hence $\mathbb{E}\{\cos{\Psi_{l,m}}\}=1$. For the case with $l\neq m$, the phase $\Psi_{l,m}$ follows the uniform random distribution in $0,2\pi)$ due to the uniformly random phases $\Phi_{l}$ and $\Phi_{m}$, and hence $\mathbb{E}\{\cos{\Psi_{l,m}}\}=0$. Therefore, the mean RSRP is given by
$\mathbb{E}\{P_{\text{r}}\}=\textstyle\sum\nolimits_{l\in\mathcal{L}}G(\Pi_{\text{tx},l})\alpha_{l}^{2}\left|\Delta_{l}\right|^{2}$ (17)
$=\textstyle\sum\nolimits_{l\in\mathcal{L}}G(\Pi_{\text{tx},l})\alpha_{l}^{2}\left|\boldsymbol{a}_{\text{tx}}^{T}(\Pi_{\text{tx},l})\,\mathbf{w}_{\text{tx}}\right|^{2}$
$=\textstyle\sum\nolimits_{l\in\mathcal{L}}G(\Pi_{\text{tx},l})\alpha_{l}^{2}\left|\boldsymbol{a}_{\text{tx,y}}^{T}(\Pi_{\text{tx},l})\,\mathbf{w}_{\text{tx,y}}\right|^{2}\left|\boldsymbol{a}_{\text{tx,z}}^{T}(\Pi_{\text{tx},l})\,\mathbf{w}_{\text{tx,z}}\right|^{2}/N_{\text{tx}}.$
Based on (3) and (8), the $y\text{-}$component of the array factor is
$\boldsymbol{a}_{\text{tx,y}}^{T}(\Pi_{\text{tx},l})\,\boldsymbol{w}_{\text{tx,y}}(\xi_{\text{tx,y}})=S_{N_{\text{tx,y}}}(\zeta_{\text{tx,y},l}+\xi_{\text{tx,y}}),$ (18)
where $\zeta_{\text{tx,y},l}\triangleq\frac{2\pi f_{\text{c}}d_{\text{y}}}{c}\mathbf{h}^{T}\boldsymbol{u}_{\text{tx}}(\Pi_{\text{tx},l})$ and $S_{N}(\cdot)$ is defined as
$S_{N}(\psi)\triangleq\textstyle\sum_{n=0}^{N-1}e^{jn\psi}=\frac{\sin(N\psi/2)}{\sin(\psi/2)}e^{j(N-1)\tfrac{\psi}{2}}.$ (19)
Similarly, the $z\text{-}$component of the array factor is given by
$\boldsymbol{a}_{\text{tx,z}}^{T}(\Pi_{\text{tx},l})\,\boldsymbol{w}_{\text{tx,z}}(\xi_{\text{tx,z}})=S_{N_{\text{tx,z}}}(\zeta_{\text{tx,z},l}+\xi_{\text{tx,z}}),$ (20)
where $\zeta_{\text{tx,z},l}\triangleq\frac{2\pi f_{\text{c}}d_{\text{z}}}{c}\mathbf{v}^{T}\boldsymbol{u}_{\text{tx}}(\Pi_{\text{tx},l})$. Therefore, we have
$\mathbb{E}\{P_{\text{r}}\}=\textstyle\sum\nolimits_{l\in\mathcal{L}}G(\Pi_{\text{tx},l})p_{l}|\Delta_{l}|^{2}=\textstyle\sum\nolimits_{l\in\mathcal{L}}\gamma_{l},$ (21)
where $\gamma_{l}\triangleq G(\Pi_{\text{tx},l})p_{l}|\Delta_{l}|^{2}$ denotes the average power gain, and $|\Delta_{l}|^{2}=\frac{1}{N_{\text{tx}}}\big{|}S_{N_{\text{tx,y}}}(\zeta_{\text{tx,y},l}+\xi_{\text{tx,y}})\big{|}^{2}\big{|}S_{N_{\text{tx,z}}}(\zeta_{\text{tx,z},l}+\xi_{\text{tx,z}})\big{|}^{2}$ denotes the power gain of the array factor $\Delta_{l}$, for path $l\in\mathcal{L}$.
Similarly, the variance of $P_{\text{r}}$ is given by
$\text{Var}\{P_{\text{r}}\}=\mathbb{E}\big{\{}(P_{\text{r}}-\mathbb{E}\{P_{\text{r}}\})^{2}\big{\}}$
$=\mathbb{E}\big{\{}\big{(}\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}\setminus\{l\}}P_{l,m}\cos{\Psi_{l,m}}\big{)}^{2}\big{\}}$
$\stackrel{{\scriptstyle(a)}}{{=}}\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}\setminus\{l\}}2P_{l,m}^{2}\mathbb{E}\{(\cos{\Psi_{l,m}})^{2}\}$
$\stackrel{{\scriptstyle(b)}}{{=}}\textstyle\sum\nolimits_{l\in\mathcal{L}}\textstyle\sum\nolimits_{m\in\mathcal{L}\setminus\{l\}}\gamma_{l}\gamma_{m}=\left(\textstyle\sum\nolimits_{l\in\mathcal{L}}\gamma_{l}\right)^{2}-\textstyle\sum\nolimits_{l\in\mathcal{L}}\gamma_{l}^{2},$ (22)
where the equation $(a)$ is due to $P_{l,m}\cos{\Psi_{l,m}}=P_{m,l}\cos{\Psi_{m,l}}$ and $\mathbb{E}\{\cos{\Psi_{l,m}}\}=0$ for $l\neq m$, and $(b)$ is due to $P_{l,m}^{2}=\gamma_{l}\gamma_{m}$ and $\mathbb{E}\{(\cos{\Psi_{l,m}})^{2}\}=1/2$. Proposition 1 thus follows. $\blacksquare$